OpenStudy (anonymous):

Which of the sums below can be expressed as 7(4 + 9)? 28 + 9 28 + 63 7 + 63 11 + 16

OpenStudy (anonymous):

i feel dumb

OpenStudy (michele_laino):

hint: \[\Large 7 \times \left( {4 + 9} \right) = 7 \times 4 + 7 \times 9 = ...?\]

OpenStudy (michele_laino):

we have to apply the distributive property of multiplication over addition

OpenStudy (michele_laino):

hint: what is: 7*4=...?

OpenStudy (michele_laino):

and what is: 7*9=...?